OpenStudy (anonymous):

Can someone please check my work?? Wanna make sure I'm doing this right! Simply each number by using the imaginary number i. thanks!

OpenStudy (anonymous):

\[\sqrt{-7}\] = 7i or \[i \sqrt{7}\]

OpenStudy (anonymous):

\[\sqrt{-81}\] = 3i

OpenStudy (anonymous):

\[\sqrt{-16}\] = \[2i \sqrt{4}\]

OpenStudy (anonymous):

\[3\sqrt{-9}\] = \[3i\]

OpenStudy (anonymous):

The last one is wrong.

OpenStudy (anonymous):

@dido525 Oh wow!! Thank you. I can't believe I'm doing these somewhat right lol. Would it be \[3i \sqrt{3}\] ?? And there was also a problem that was \[\sqrt{-72}\] I got: \[2i \sqrt{3}\]

OpenStudy (anonymous):

The last one should be 9i.

OpenStudy (anonymous):

Well think about it WITHOUT the negative.

OpenStudy (anonymous):